OpenStudy (anonymous):

The polygons below are similar. Find the value of y. 4.5 7.5 12 16

OpenStudy (anonymous):

i think its a 4.5 but can u explain how u get the answer

OpenStudy (anonymous):

Well, i'll set up a proportion, \[\frac{ 8 }{ 6 } = \frac{ 6 }{ }\]

OpenStudy (sissyedgar):

i think so 2, but i cant explain cause i kinda guessed

OpenStudy (anonymous):

8/6 = 6/y

OpenStudy (anonymous):

So 8y=36, solve y for your answer

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Well its like one shape over the other

OpenStudy (anonymous):

like the way ur putting it doesnt make sense like can u walk me thru each step

OpenStudy (anonymous):

I had BC over FG equals AB over EF

OpenStudy (anonymous):

You're trying to set up a proportion with the shapes

jimthompson5910 (jim_thompson5910):

this may help

OpenStudy (anonymous):

im bad a proporations when y is in it

OpenStudy (anonymous):

That image posted is perfect for explaining. (I'm bad at it) xD

OpenStudy (anonymous):

still dont understand one bit

jimthompson5910 (jim_thompson5910):

There are a few ways to set up the problem. One of them is this proportion 6/y = 8/6 cross multiply and solve for y to get 6*6 = y*8 36 = 8y 8y = 36 y = ???

OpenStudy (anonymous):

4.5

OpenStudy (sissyedgar):

I get it!

OpenStudy (anonymous):

\[\frac{ BC }{ FG } = \frac{ AB }{ EF } use the imagetohelp\]